OpenStudy (anonymous):

Help Me!!!! 6 = 2(x + 8) - 5x

OpenStudy (skullpatrol):

Do you want to solve for x?

OpenStudy (anonymous):

Solve equation The options none of the answers given

OpenStudy (skullpatrol):

You made a mistake brotha...

OpenStudy (phoenixfire):

@bunyonb You missed something... it's 5x. not 5 in the start.

OpenStudy (skullpatrol):

That is why we let the asker solve BROTHA

OpenStudy (skullpatrol):

pwned

OpenStudy (phoenixfire):

6 = 2(x + 8) - 5x 1) Expand brackets 2) Rearrange so all terms with an x are on one side 3) combine like terms, and solve.

OpenStudy (anonymous):

A. 2/3 B. 3 1/3 C. - 2/3 D. -3 1/3

OpenStudy (phoenixfire):

@bunyonb Stop! Let the person who asked learn by doing it.

OpenStudy (phoenixfire):

@Tabbiejack Try it. Read what I said, and type out your steps.

OpenStudy (skullpatrol):

later brotha

OpenStudy (anonymous):

I am lost bad

OpenStudy (phoenixfire):

Start by expanding the brackets.

OpenStudy (skullpatrol):

Ask yourself " What do you want to solve for?"

OpenStudy (anonymous):

Ok

OpenStudy (skullpatrol):

"Solve for" means you want to get that letter alone on one side of the equals sign, right?

OpenStudy (anonymous):

10/3x

OpenStudy (anonymous):

But that's not an option

OpenStudy (skullpatrol):

What happened to the equals sign?

OpenStudy (anonymous):

6=2x+16-5x 6=-3x+16 6-16=-3x -10=-3x -10/-3=x 10/3=x

OpenStudy (anonymous):

I'm lost

OpenStudy (anonymous):

People I'm so freaking lost imabout to cry

OpenStudy (skullpatrol):

First thing is calm down, please :-D

OpenStudy (skullpatrol):

Check your answer in the original equation

OpenStudy (phoenixfire):

10/3 is correct. what you need to do is simplify it. how many 3s fit into 10? three, right? with 1 left. So your simplified one is \[{10 \over 3}=>3{1\over 3}\]

OpenStudy (anonymous):

O ooh I feel so stupid

OpenStudy (skullpatrol):

6 = 2(x + 8) - 5x <----original equation

OpenStudy (anonymous):

Just 1 more thing. - 1/3m - 7 = 5

OpenStudy (help!!!!):

3 1/3

OpenStudy (anonymous):

I got it thank its the other 1